OpenStudy (anonymous):

Which expression is a fourth root of -1+isqrt3?

OpenStudy (anonymous):

4th root(cos(60) + isin(60)) 4th root(cos(30) + isin(30)) 4th root(cos(280) + isin(280)) 4th root(cos(320) + isin(320))

OpenStudy (anonymous):

One way of looking at this is to find the \(\theta\) that satisfies the following.\[\sqrt[4]{-1+i \sqrt 3}=\sqrt[4]{\cos \theta + i \sin \theta}\]Just by looking at the above equation, what can we say that \(\cos \theta\) and \(\sin \theta\) equal?

OpenStudy (anonymous):

Im not quite sure, can you please just provide the answer? So i can figure it out myself

OpenStudy (anonymous):